# Interpretable Multimodal Learning for Tumor Protein-Metal Binding: Progress, Challenges, and Perspectives

### v1

@@ 6 Perspectives and future directions
In this section, we summarized key challenges that hinder progress in interpretable multimodal learning for tumor protein–metal binding. We then discussed potential solutions based on insights from previous sections. Beyond these solutions, we highlighted two emerging research directions that extend current efforts and have the potential to advance this domain further.
6.1 Challenges and perspectives for tumor protein-metal binding prediction
@@
Several technical and practical barriers remain unresolved in this domain. First, the scarcity of high-quality, tumor-specific datasets limits the training and generalization capabilities of ML models. Second, integrating diverse data modalities, such as protein sequences, structures, binding pockets, and textual annotations, introduces complexity, as each carries distinct formats and levels of biological insights. Developing robust fusion strategies that preserve meaningful signals across modalities remains an open problem. Finally, the “black-box” nature of complex ML models complicates the interpretation of their decision-making processes. This lack of transparency is particularly concerning in biomedical research, where model interpretability is critical for ensuring scientific credibility and facilitating clinical translation.
To address these challenges, we discussed the following perspectives on potential solutions:
  1. 1.
@@
    Integrating cancer-related and protein metal-binding datasets: Combining existing tumor protein datasets (e.g., TCGA) with metal-binding datasets (e.g., PDB, MetalPDB) can help address data scarcity and generate more diverse, informative training datasets for ML model development.
  2. 2.
@@
    Adapting existing models for tumor-specific prediction: General protein–metal binding models (e.g., LMetalSite, M-Ionic) can be retrained on tumor-specific data to support targeted prediction tasks. Additionally, pre-trained protein foundation models can be fine-tuned for tumor-relevant applications.
  3. 3.
@@
    Leveraging multimodal learning with structured data workflows: Implementing a well-defined data processing and feature extraction workflow ensures high-quality inputs for ML models. Pairing structured workflows with appropriate multimodal fusion strategies and ML models can maximize the complementary strengths of diverse data modalities, leading to improved model performance and robustness.
  4. 4.
@@
    Improving model interpretability: Incorporating inherently interpretable models (e.g., decision trees, linear models) and applying post-hoc explanation techniques (e.g., LIME, SHAP) can help uncover biologically meaningful insights and promote trust in ML-driven predictions.
@@
In addition to the proposed solutions, we identified two emerging directions that promise to advance tumor protein–metal binding prediction by deepening biological insight and expanding modeling capabilities. In particular, integrating protein–protein interaction (PPI) data can provide structural context for metal-binding events, while modeling post-binding structure changes in tumor proteins offers new insight into how metal coordination alters protein function. These directions represent promising extensions of current efforts, with the potential to enhance both mechanistic understanding and therapeutic development.
6.2 Integrating protein–protein interactions into tumor protein–metal binding prediction
@@
Protein-protein interaction (PPI) networks are essential in cellular function, and tumor proteins often exert their effects through these networks. Incorporating PPI information can provide valuable functional context for understanding the biological significance and functional impact of metal-binding events. For example, the organoarsenic drug Darinaparsin (ZIO-101) has demonstrated anticancer activity across various tumor cell lines. Studies has shown that ZIO-101 can bind to histone H3.3, affecting the function of histone H3.3 in the cell nucleus [142]. PPI analysis reveals that histone H3.3 closely interacts with histone deacetylase 1 (HDAC1). HDAC1 can regulate the level of histone acetylation, a process that further affects the expression of cyclin-dependent kinase inhibitor 1A (CDKN1A), ultimately leading to the overexpression of tumor necrosis factor-related genes and triggering apoptosis, thus inhibiting tumor growth. Focusing solely on the binding of histone H3.3 to arsenic without considering its interaction with HDAC1 would fail to fully grasp the anticancer mechanism of arsenic compounds and could overlook potential therapeutic strategies targeting HDAC1 [142]. These examples highlight that PPI often influence the formation, stability, and function of metal-binding sites. PPI data can enhance model accuracy by providing biologically relevant context [98]. However, current prediction models make limited use of PPI information. Therefore, we encourage future research to integrate PPI networks to improve the functional relevance and interpretability of tumor protein–metal binding predictions.
@@
Recent advances in bioinformatics have provided a solid foundation for this integration. Public datasets, such as STRING [143] and PINA [144], have accumulated extensive PPI data that can be mined to extract interaction networks relevant to tumor protein-metal binding functions. Meanwhile, computational methods such as molecular docking and molecular dynamics simulation allow researchers to model protein-protein interactions and complexes, and examine how metal-binding sites are altered in these contexts [145]. In addition, experimental techniques such as yeast two-hybrid screening and co-immunoprecipitation can validate predicted interactions, providing strong experimental support for ML models [146, 147]. User-friendly PPI network analysis platforms improve accessibility for researchers with limited programming expertise, facilitating broader adoption of PPI-informed modeling strategies [148].
@@
6.3 Post-binding structural prediction in tumor protein-metal complexes
@@
The 3D structure of a protein determines its cellular function, including catalysis, signal transduction, and structural support. When binding to a metal ion, proteins often undergo structural changes that alter their function. This also holds for tumor proteins, where metal binding may disrupt or modulate oncogenic processes. Metal-based drugs, such as platinum-based drugs, exploit this mechanism to alter tumor protein structures and induce cell death.
To model this process computationally, we propose leveraging existing protein–metal binding predictive methods and extend them to estimate structural changes in tumor proteins after metal binding. These models can serve as valuable tools for assessing therapeutic potential, accelerating drug development, and overcoming the limitations of experimental structural determination. Existing models for studying protein-metal interactions operate within a ternary framework composed of three key components: the protein, its binding site(s), and the associated metal ions. For example, models using protein and binding sites data as inputs can predict the types of metal that may bind. Conversely, models using protein and metal information can identify whether a protein functions as a metalloprotein and locate its potential binding sites. Through these modular configurations, models can infer conformational changes even with incomplete input data, enabling structure prediction under conditions where structural data are missing or limited [98].
Based on the insights from recent studies [149, 84], we propose a framework for post-binding structural prediction in tumor protein-metal complexes, as shown in Fig. 3. The framework operates in two stages: first, it predicts binding sites using tumor protein and metal ion inputs; then, it uses the protein, metal, and predicted binding sites to model the post-binding structure. To enhance prediction accuracy, the framework incorporates prior knowledge from known metalloprotein structures and metal-specific structural effects. Additionally, domain-specific constraints, such as chirality and geometric coordination rules, are applied to ensure biological plausibility. This framework supports the predictive modeling of metal-bound tumor protein structures and offers a pathway toward structure-guided drug design targeting metal–protein interactions in cancer.
@@ 7 Conclusion
Interpretable multimodal learning offers a promising approach to improving the efficiency of wet-lab experiments in tumor protein-metal binding research while ensuring model transparency and predictive accuracy. In this paper, we presented our perspective on advancing this field by drawing insights from existing protein-metal binding studies among data, multimodal learning, and interpretability. We highlighted data modalities and key data processing steps for conducting tumor protein-metal binding research, including standardizing data acquisition, optimizing multimodal learning, and improving interpretability. Building on these insights, we proposed two additional considerations for advancing future research in tumor protein-metal binding: integrating protein-protein interaction data and developing models capable of predicting structural changes in tumor proteins after metal binding. We believe that implementing these perspectives will establish a strong foundation for future interdisciplinary research and will require close collaboration among clinicians, biologists, and machine learning researchers.

### v2

@@ 6 Perspectives and future directions
Finally, we summarize key challenges that hinder progress in interpretable multimodal learning for tumor protein–metal binding and discuss promising directions based on insights from previous sections. Beyond these solutions, we highlight two emerging research directions that extend current efforts and have the potential to further advance this domain.
6.1 Challenges and opportunities in prediction
@@
Several technical and practical barriers remain unresolved in this domain. First, the scarcity of high-quality, tumor-specific datasets limits the training and generalization capabilities of ML models. Second, integrating diverse data modalities, such as protein sequences, structures, binding pockets, and textual annotations, introduces complexity, as each carries distinct formats and levels of biological insight. Developing robust fusion strategies that preserve meaningful signals across modalities remains a significant challenge. Lastly, the “black-box” nature of complex ML models remains a barrier to transparency, particularly in biomedical research where interpretability is essential for credibility and clinical translation.
To address these challenges, we explore the following promising directions:
@@
  • Integrating cancer-related and protein-metal binding datasets: Combining existing tumor protein datasets (e.g., TCGA) with protein-metal binding datasets (e.g., PDB, MetalPDB) can help address data scarcity and generate more diverse, informative training datasets for ML model development.
@@
  • Adapting existing models for tumor-specific prediction: Existing protein–metal binding models (e.g., LMetalSite, M-Ionic) can be adapted to tumor-specific settings to support targeted prediction tasks. Additionally, pre-trained protein foundation models can be fine-tuned and integrated for tumor-relevant applications.
@@
  • Leveraging multimodal learning with standardized data workflows: Standardizing well-defined data processing and feature extraction workflows ensures high-quality inputs for ML models. Pairing standardized workflows with appropriate multimodal fusion strategies and ML models can maximize the complementary strengths of diverse data modalities, leading to improved model performance and robustness.
@@
  • Improving model interpretability: Incorporating inherently interpretable ML models (e.g., decision trees, linear models) and other by-design interpretable ML models and applying post hoc explanation methods (e.g., LIME, SHAP), while addressing their limitations, can help uncover biologically meaningful insights and promote trust in ML-driven predictions.
Together, these strategies provide a foundation for improving prediction accuracy, transparency, and translational potential. In addition to these strategies, biological complexity presents further challenges for accurately predicting tumor protein–metal binding. The tumor microenvironment is a dynamic and heterogeneous system in which factors such as metal ion concentrations and oxidative stress can influence both the likelihood and outcome of metal-binding events. One way to address this complexity is through multi-omics integration models that incorporate metabolomics-derived features, such as metal ion concentrations and oxidative stress biomarkers, together with protein structural data to help elucidate the dynamic distribution of metals within the tumor microenvironment and their impact on protein conformation [152, 153]. Incorporating such context-aware biological information may support more realistic modeling of metal–protein interactions in vivo and motivate future integration of spatiotemporal data and simulation-based approaches, ultimately bridging the gap between predictive modeling and real-world biological systems.
@@
Beyond these practical strategies, we highlight two emerging directions that could transform our understanding of tumor protein–metal interactions: integrating protein–protein interaction (PPI) data can provide structural context for metal-binding events, while modeling post-binding structure changes in tumor proteins offers new insights into how metal coordination reshapes structure and function. These directions represent promising extensions of current efforts, with the potential to enhance both mechanistic understanding and therapeutic development.
6.2 Integrating protein–protein interaction networks for contextual prediction
@@
Protein-protein interaction (PPI) networks are essential in cellular function, and tumor proteins often exert their effects through these networks. Incorporating PPI information can provide valuable functional context for understanding the biological significance and functional impact of metal-binding events. For example, the organoarsenic drug Darinaparsin (ZIO-101) has demonstrated anticancer activity across various tumor cell lines. Studies have shown that ZIO-101 can bind to histone H3.3, affecting the function of histone H3.3 in the cell nucleus [154]. PPI analysis reveals that histone H3.3 closely interacts with histone deacetylase 1 (HDAC1). HDAC1 can regulate the level of histone acetylation. This process further affects the expression of cyclin-dependent kinase inhibitor 1A (CDKN1A). Ultimately, it leads to the overexpression of genes related to tumor necrosis factor, which induces apoptosis and thus inhibits tumor growth. Focusing solely on the binding of histone H3.3 to arsenic without considering its interaction with HDAC1 would fail to fully grasp the anticancer mechanism of arsenic compounds and could overlook potential therapeutic strategies targeting HDAC1 [154]. These examples highlight that PPI often influence the formation, stability, and function of metal-binding sites. PPI data can enhance model accuracy by providing biologically relevant context [100]. However, current prediction models make limited use of PPI information. Therefore, we encourage future research to integrate PPI networks to improve the functional relevance and interpretability of tumor protein–metal binding predictions.
@@
Recent advances in bioinformatics have provided a solid foundation for this integration. Public datasets, such as STRING [155] and PINA [156], have accumulated extensive PPI data that can be mined to extract interaction networks relevant to tumor protein-metal binding functions. At the same time, computational methods such as molecular docking and molecular dynamics simulations enable researchers to model protein-protein interactions and complexes. They can also investigate how metal binding sites change in these cases [157]. In addition, experimental techniques such as yeast two-hybrid screening and coimmunoprecipitation can validate predicted interactions, providing strong experimental support for ML models [158, 159]. User-friendly PPI network analysis platforms improve accessibility for researchers with limited programming expertise, facilitating broader adoption of PPI-informed modeling strategies [160]. As such, integrating PPI data can significantly enhance the biological fidelity and interpretability of tumor metal-binding predictions.
@@
6.3 Modeling post-binding structural changes in tumor protein–metal complexes
The 3D conformation of a protein underpins its biological function, including catalysis, signal transduction, and structural support. When binding to a metal ion, proteins often undergo structural changes that alter their function. This also holds for tumor proteins, where metal binding may disrupt or modulate oncogenic processes. Metal-based drugs, such as platinum-based drugs, exploit this mechanism to alter tumor protein structures and induce cell death.
To model this process computationally, we propose using existing protein–metal binding predictive methods and extending them to estimate the structural changes in tumor proteins after metal binding. These models can serve as valuable tools for assessing therapeutic potential, accelerating drug development, and overcoming the limitations of experimental structural determination. Existing models for studying protein-metal interactions operate within a ternary framework composed of three key components: the protein, its binding site(s), and the associated metal ions. For example, models using protein and binding site data as inputs can predict the types of metals that may bind. Conversely, models using protein and metal information can identify whether a protein functions as a metalloprotein and locate its potential binding sites. Through these modular configurations, models can infer conformational changes even with incomplete input data, enabling structure prediction under conditions where structural data are missing or limited [100].
Based on recent studies [161, 86], we propose a framework for post-binding structural prediction in tumor protein-metal complexes, as shown in Fig. 3. The framework operates in two stages: first, it predicts binding sites using inputs from tumor proteins and metal ions; then, it utilizes the protein, metal, and predicted binding sites to model the post-binding structure. To enhance prediction accuracy, the framework incorporates prior knowledge from known metalloprotein structures and metal-specific structural effects. Additionally, domain-specific constraints, such as chirality and geometric coordination rules, are applied to ensure biological plausibility. This framework facilitates the predictive modeling of metal-bound tumor protein structures, providing a pathway toward structure-guided drug design targeting metal–protein interactions in cancer.
@@ 7 Conclusion
Interpretable multimodal learning presents a promising avenue for enhancing the efficiency and transparency of tumor protein–metal binding research. In this paper, we have outlined a perspective shaped by insights from existing protein–metal binding studies, encompassing diverse data modalities, multimodal learning strategies, and model interpretability. We highlighted the importance of standardizing data acquisition, optimizing multimodal integration, and developing interpretable machine learning models. Beyond these core areas, we proposed two emerging directions for advancing future research: integrating protein–protein interaction data to provide functional context, and modeling structural changes in tumor proteins following metal binding to inform therapeutic development. We believe that realizing this vision will require close collaboration between clinicians, biologists, and machine learning researchers, establishing a strong foundation for impactful and interdisciplinary progress in the field.